On a map, two cities are 3 ¾ inches apart. If 3/8 inches on the map represents 50 miles, find the number of miles between the cities.

Answer by mananth(16949)

3/8 in = 50 miles
...
3 3/4 = 15/4
...
15/4 will represent ?
..
direct proportion
((15/4) *50)/(3/8)
(325/2)/(3/8)
(325/2)*(8/3)
=433.33 miles
...
m.ananth@hotmail.ca